OpenStudy (anonymous):

an air force plane flew to istanbul and back. the trip there took 5 hours and the trip back took 4 hours. what was the air force plane's average speed on the trip there if it averaged 240 mph on the return trip?

OpenStudy (anonymous):

Trip back: S= 240mph, T= 4h Dist= s X T = 240 X 4 =960 m

OpenStudy (anonymous):

Av speed = total dist/total time taken =(960 x 2)/ (5+4) = 213 1/3 mph

OpenStudy (anonymous):

Sorry....av speed to istanbul= Dist / Time =960/5 = 192 mph

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

no prob